OpenStudy (anonymous):

step 1: Solve quadratic equation x2−8x+12=0 by using step-by-step quadratic equation solver (opens new window). The solutions are: x1=2 x2=6 step2: Use formula for factoring quadratic equation: @ganeshie8 @Nnesha @iGreen. @confluxepic

OpenStudy (anonymous):

oops

Nnesha (nnesha):

\[\huge\rm x = \frac{ -b \pm \sqrt{b2 -4ac} }{ 2a }\]

OpenStudy (anonymous):

2. Heinz has a list of possible functions. Pick one of the g(x) functions below, show how to find the zeros, and then describe to Heinz the other key features of g(x). • g(x) = x3 – x2 – 4x + 4 • g(x) = x3 + 2x2 – 9x – 18 • g(x) = x3 – 3x2 – 4x + 12 • g(x) = x3 + 2x2 – 25x – 50 • g(x) = 2x3 + 14x2 – 2x – 14 3. Provide a rough sketch of g(x). Label or identify the key features on the graph

OpenStudy (anonymous):

thats my question and btw its a study guide not a graded assignement

Nnesha (nnesha):

do you know about group method ??

OpenStudy (anonymous):

not at all.

Nnesha (nnesha):

 i'm doing 3rd one make 2 group ^^